# Single-Shot Readout and Weak Measurement of a Tin-Vacancy Qubit in Diamond

### v1

@@ A.1.3 Comparison between models
To confirm these models, we compare simulations of the three-level model in Sec. A.1.1 to the two-level model in Sec. A.1.2. In each case, we numerically solve the master equation using the QuTiP package [61].
Figure 7: (a) Dynamics of a driven 3-level atomic system, Fig. 6, simulated using the Hamiltonian and collapse operators given by Eq. 11-13. (b) Dynamics of the simplified 2-level model, simulated using collapse operators Eq. 14 and Eq. 15. Both (a) and (b) use $\chi=100$, $\gamma/2\pi=35\,\mathrm{MHz}$, $\Omega_{\mathrm{A1}}=5\gamma$, and $\delta=0$. (c) Polarization rate $\kappa$ (left) and dephasing rate $\Gamma_{\phi}$ (right) as functions of $\delta$, and at different values of $\Omega_{\mathrm{A1}}/\gamma$. Numerical solutions to the 3-level model described in Appendix A.1.1 (dots) match the analytical solutions (Eq. 17-Eq. 20) based on the 2-level model in Appendix A.1.2 (lines).
The time-domain dynamics of both models are plotted in Fig. 7. In both cases the system undergoes spin polarization at rate $\kappa$ and an overall envelope of drive-induced dephasing at rate $\Gamma_{\phi}$. The three-level model also captures the coherent dynamics between the spin ground states and the |$\ket{A}$ state. We simulate both models as a function of detuning $\delta$, and confirm they match. This comparison also illustrates power broadening – the effective increase in linewidth at high drive powers.
@@ A.2 Emission model
A driven atomic system emits at the steady-state photon scattering rate given by Eq. 16, times the population in the driven spin state [41]. For example, upon onset of a laser drive at time $t=0$ the emission rate is:
$\Gamma_{\mathrm{em}}(t)=|\alpha|^{2}Re^{-\kappa t},$ (21)
@@
In readout, “dark” or “bright” distributions with low ($\bar{n}_{d}$) or high ($\bar{n}_{b}$) mean photon numbers are associated with the qubit being prepared in different spin states prior to measurement. For example, state preparation of $\ket{\uparrow}$ yields $\alpha=0$, no initial emission, and a dark collected distribution. State preparation of $\ket{\downarrow}$ yields $\alpha=1$, maximum initial emission, and a bright collected distribution. However, emitted photons are only collected with a probability $\eta$ (the measurement efficiency) and furthermore, collected counts can also come from unwanted sources unrelated to the qubit, i.e. noise arriving at rate $\Gamma_{\mathrm{noise}}$. We therefore model $\bar{n}_{d}$ and $\bar{n}_{b}$ as:
@@
where $\tau$ is the readout integration window. We model noise to be time-independent arriving at rate $\Gamma_{\mathrm{noise}}=a+bp$, where $a$ is the mean rate of background noise photons including dark counts, $b$ is noise from scatter of the excitation laser into the collection path and is proportional to power. We fit this data to Eq. 22 to determine $a\tau=0.028\pm 0.002\,\mathrm{photons}$ and $b\tau=(6.24\pm 0.15)\times 10^{-5}\,\mathrm{photons/nW}$.
For preparation in the spin state associated with maximum emission ($\alpha=1$), the difference in collected counts between the bright and dark distributions is therefore: $\bar{n}_{b}-\bar{n}_{d}=\eta\chi(1-e^{-\kappa\tau})$. For $\kappa\tau\gg 1$, so the duration of drive is long enough to entirely polarize the spin state, this reduces to $\bar{n}_{b}-\bar{n}_{d}=\eta(\chi$+1). In other words, the expected number of detection events is essentially the cyclicity $\chi$ times the measurement efficiency $\eta$.
A.3 Readout fidelity vs. power
We now use this model of emission to study readout fidelity as a function of power. Readout fidelity quantifies the distinction between the “bright” or “dark” distributions of photons collected when the qubit is prepared in the $\ket{\downarrow}$ or $\ket{\uparrow}$ spin state. These distributions are distinguished by a readout threshold $N_{r}$, meaning the outcome “0” can be associated with the collection of $\geq N_{r}$ counts per readout window, and the outcome “1” with $<N_{r}$ counts. In practice, $N_{r}$ is chosen to maximize readout fidelity $F_{r}=1-\frac{1}{2}P(1|\downarrow)-\frac{1}{2}P(0|\uparrow)=\frac{1}{2}P(0|\downarrow)+\frac{1}{2}P(1|\uparrow)$.
Here, $P(1|\downarrow)$ and $P(0|\uparrow)$ are error probabilities, associated with the chance a collected number of counts falls within a different distribution than expected given state preparation. Infidelity can result from too many counts in the dark distribution or too few counts in the bright distribution. It can also result from state preparation errors $\epsilon_{1}$ ($\epsilon_{0}$); the chances the qubit was actually prepared in the $\ket{\downarrow}$ ($\ket{\uparrow}$) state when it was attempted to be prepared in the $\ket{\uparrow}$ ($\ket{\downarrow}$) state, respectively.
To model readout fidelity we assume that emission follows Poissonian statistics arising from a Markovian process. These statistics are known to describe emission from atomic systems [62], but can be modified in other experiments where interaction with a photonic structure changes the electromagnetic density of states [33]. Poissoinian statistics are captured by the model in Eq. A.1, which leads to the emission rate in Eq. 21. Therefore during a readout event, collection of $k$ counts is expected with probability:
@@
The error probabilities $P(1|\downarrow)$ and $P(0|\uparrow)$ are thus:
@@
$P(0|\uparrow)=\sum_{k=N_{r}}^{\infty}\epsilon_{1}p[k,\bar{n}_{b}]+(1-\epsilon_{1})p[k,\bar{n}_{d}].$ (26)
Evaluating using the statistics given by Eq. 24 yields:

### v2

@@ A.1.3 Comparison between models
To confirm these models, we compare simulations of the three-level model in Appendix A.1.1 to the two-level model in Appendix A.1.2. In each case, we numerically solve the master equation using the QuTiP package [70].
Figure 7: (a) Dynamics of a driven three-level atomic system, Fig. 6, simulated using the Hamiltonian and collapse operators given by Eqs. 11-13. (b) Dynamics of the simplified two-level model, simulated using collapse operators Eqs. 14 and 15. Both (a) and (b) use $\Lambda=100$, $\gamma/2\pi=35\,\mathrm{MHz}$, $\Omega_{\mathrm{A1}}=5\gamma$, and $\delta=0$. (c) Polarization rate $\Gamma_{p}$ (left) and dephasing rate $\Gamma_{\phi}$ (right) as functions of $\delta$ and at different values of $\Omega_{\mathrm{A1}}/\gamma$. Numerical solutions to the three-level model described in Appendix A.1.1 (dots) match the analytical solutions (Eqs. 17-20) based on the two-level model in Appendix A.1.2 (lines).
The time-domain dynamics of both models are plotted in Fig. 7. In both cases, the system undergoes spin polarization at rate $\Gamma_{p}$ and an overall envelope of drive-induced dephasing at rate $\Gamma_{\phi}$. The three-level model also captures the coherent dynamics between the spin ground states and the |$\ket{A}$ state. We simulate both models as a function of detuning $\delta$ and confirm they match. This comparison also illustrates power broadening—the effective increase in linewidth at high drive powers.
@@ A.2 Emission model
A driven atomic system emits at the steady-state photon scattering rate given by Eq. 16, times the population in the driven spin state [45]. For example, upon onset of a laser drive at time $t=0$, the emission rate is
$\Gamma_{\mathrm{em}}(t)=|\alpha|^{2}Re^{-\Gamma_{p}t},$ (21)
@@
In readout, “dark” or “bright” distributions with low ($\bar{n}_{d}$) or high ($\bar{n}_{b}$) mean photon numbers are associated with the qubit being prepared in different spin states prior to measurement. For example, state preparation of $\ket{\uparrow}$ yields $\alpha=0$, no initial emission, and a dark collected distribution. State preparation of $\ket{\downarrow}$ yields $\alpha=1$, maximum initial emission, and a bright collected distribution. However, emitted photons are collected only with a probability $\eta$ (the measurement efficiency), and, furthermore, collected counts can also come from unwanted sources unrelated to the qubit, i.e., noise arriving at rate $\Gamma_{\mathrm{noise}}$. We, therefore, model $\bar{n}_{d}$ and $\bar{n}_{b}$ as
@@
respectively, where $\tau$ is the readout integration window. We model noise to be time independent arriving at rate $\Gamma_{\mathrm{noise}}=a+bp$, where $a$ is the mean rate of background noise photons including dark counts, $b$ is noise from scatter of the excitation laser into the collection path and is proportional to power.
For preparation in the spin state associated with maximum emission ($\alpha=1$), the difference in collected counts between the bright and dark distributions is, therefore, $\bar{n}_{b}-\bar{n}_{d}=\eta(\Lambda+1)(1-e^{-\Gamma_{p}\tau})$. In Eq. 6, we simplify this by taking $(\Lambda+1)\to\Lambda$, since in this work we operate in the $\Lambda\gg 1$ limit where $(\Lambda+1)\approx\Lambda$. For $\Lambda\tau\gg 1$, so the duration of drive is long enough to entirely polarize the spin state, this reduces to $\bar{n}_{b}-\bar{n}_{d}\approx\eta\Lambda$. In other words, the expected number of detection events is essentially the cyclicity $\Lambda$ times the measurement efficiency $\eta$.
A.3 Readout fidelity vs power
We now use this model of emission to study readout fidelity as a function of power. Readout fidelity quantifies the distinction between the bright or dark distributions of photons collected when the qubit is prepared in the $\ket{\downarrow}$ or $\ket{\uparrow}$ spin state. These distributions are distinguished by a readout threshold $N_{r}$, meaning the outcome “0” can be associated with the collection of $\geq N_{r}$ counts per readout window and the outcome “1” with $<N_{r}$ counts. In practice, $N_{r}$ is chosen to maximize readout fidelity $F_{r}=1-\frac{1}{2}P(1|\downarrow)-\frac{1}{2}P(0|\uparrow)=\frac{1}{2}P(0|\downarrow)+\frac{1}{2}P(1|\uparrow)$.
Here, $P(1|\downarrow)$ and $P(0|\uparrow)$ are error probabilities, associated with the chance a collected number of counts falls within a different distribution than expected given state preparation. Infidelity can result from too many counts in the dark distribution or too few counts in the bright distribution. It can also result from state preparation errors $\epsilon_{1}$ ($\epsilon_{0}$), the chances the qubit was actually prepared in the $\ket{\downarrow}$ ($\ket{\uparrow}$) state when it was attempted to be prepared in the $\ket{\uparrow}$ ($\ket{\downarrow}$) state, respectively.
To model readout fidelity, we assume that emission follows Poissonian statistics arising from a Markovian process. These statistics are known to describe emission from atomic systems [71] but can be modified in other experiments where interaction with a photonic structure changes the electromagnetic density of states [37]. Poissonian statistics are captured by the model in Eq. A.1, which leads to the emission rate in Eq. 21. Therefore, during a readout event, collection of $k$ counts is expected with probability
@@ A.3 Readout fidelity vs power
The error probabilities $P(1|\downarrow)$ and $P(0|\uparrow)$ are, thus,
@@
$P(0|\uparrow)=\sum_{k=N_{r}}^{\infty}\epsilon_{1}p[k,\bar{n}_{b}]+(1-\epsilon_{1})p[k,\bar{n}_{d}],$ (26)
respectively. Evaluating using the statistics given by Eq. 24 yields
@@
A.4 Spin readout using cavity quantum electrodynamics
As discussed in the main text Sec. IV, comparison of emission to measurement-induced dephasing allows for a general way to characterize measurement efficiency, defined as the total loss between qubit and detector. In this section, we show how this idea is applicable to a strongly interacting cavity QED system, driven near resonance to interrogate the spin state. This style of spin readout has previously been demonstrated using SiV- centers in diamond [19, 20, 47, 22] and is a natural next step for the SnV- platform.
To understand how spin readout works in a probed cavity QED system, consider a qubit whose state is described by density matrix $\hat{\rho}$ and which is prepared in the spin superposition (|+|$(\ket{\downarrow}+\ket{\uparrow})/\sqrt{2}$. This spin may interact in a spin-selective way with another quantum state such as the electromagnetic field in a driven cavity. Upon interaction, this other state can acquire a spin-dependent phase and/or amplitude shift, such that the joint quantum state of the spin and cavity is [58, 67]
[…]
with density matrix $\hat{\sigma}=\ket{\psi}\bra{\psi}$ |. Here, |$\ket{\alpha_{\downarrow}}$ and |$\ket{\alpha_{\uparrow}}$ are the quantum states of light in the driven cavity, and are projectively measured to gain information about the spin.
Figure 8: (a) Spin readout discussed thus far in this work is based on the presence or absence of photoluminescence under a resonant optical pump. Alternatively, in a cavity quantum electrodynamics system (b) for which an SnV-’s optical transition interacts with a cavity, the spin may be measured via spin-dependent scattering from a probe tone. (c) This technique, known as “dispersive readout” requires the cavity to have a spin-dependent dispersive shift $\chi$ that is appreciable compared to the cavity linewidth $\kappa$ [20, 19, 67, 72]. (d) For both techniques, the measurement efficiency $\eta$ may be precisely determined from a comparison between measurement-induced dephasing and the rate at which information is acquired by a measurement apparatus [58]. The top left of this cartoon plot is inaccessible, as it would indicate the impossible case of a detector which gains information about the qubit state without sufficient collapse of its superposition, and $\eta=1$ bounds physically realizable measurements.
A.4.1 Measurement-induced dephasing of a cavity QED system
The overlap between |$\ket{\alpha_{\downarrow}}$ and |$\ket{\alpha_{\uparrow}}$ sets both how much the spin state may be distinguished by measurement [58, 67] and how much dephasing is caused by measurement.
To show this, following the discussion in Ref. [60] consider a general quantum measurement described by the positive operator-valued measure (POVM) $\hat{\Pi}_{n}=\hat{K}_{n}^{\dagger}\hat{K}_{n}$. Here, $\hat{K}_{n}^{\dagger}$ and $\hat{K}_{n}$ are Krauss operators that correspond to measurement of a light field in the number basis, and $n$ is a non-negative integer. After measurement, the quantum state of the joint spin-photonic system is described by density matrix $\hat{\sigma}^{\prime}=\sum_{n}\hat{K}_{n}\hat{\sigma}\hat{K}_{n}^{\dagger}$ [73].
We model readout of the joint system $\hat{\sigma}$ using the POVM $\hat{\Pi}_{n}=\mathbb{I}\otimes\ket{n}\bra{n}$ |, which describes measurement of the second system in the basis |$\ket{n}$. After a measurement outcome $n$, the joint system is:
[…]
where $\nu_{n}=\mathrm{Tr}\left(\hat{\Pi}_{n}\hat{\sigma}\right)$ is a normalization factor.
Taking the partial trace $\hat{\rho}^{{}^{\prime}}=\mathrm{Tr}_{2}\left(\sum_{n}\nu_{n}\hat{\sigma}_{n}^{{}^{\prime}}\right)$ over the second system gives the spin’s postmeasurement density matrix:
$\hat{\rho}^{{}^{\prime}}=\frac{1}{2}\begin{pmatrix}1&\braket{\alpha_{\uparrow}}{\alpha_{\downarrow}}\\ \braket{\alpha_{\downarrow}}{\alpha_{\uparrow}}&1\end{pmatrix}.$ (31)
From Eq. 31, we see that spin coherence is reduced to:
$\hat{\rho}_{\downarrow\uparrow}^{\prime}=\frac{1}{2}\braket{\alpha_{\downarrow}}{\alpha_{\uparrow}}.$ (32)
For example, consider a cavity that is in one of two spin-dependent coherent states:
|$\ket{\alpha_{\downarrow}}=e^{-\frac{1}{2}|\alpha_{\downarrow}|^{2}}\sum_{n=0}^{\infty}\frac{\alpha_{\downarrow}^{n}}{\sqrt{n!}}\ket{n},$ (33)
|$\ket{\alpha_{\uparrow}}=e^{-\frac{1}{2}|\alpha_{\uparrow}|^{2}}\sum_{n=0}^{\infty}\frac{\alpha_{\uparrow}^{n}}{\sqrt{n!}}\ket{n}.$ (34)
After the cavity state is projectively measured, the spin’s coherence is
$|\hat{\rho}_{\downarrow\uparrow}^{\prime}|=\frac{1}{2}e^{-|\alpha_{\downarrow}-\alpha_{\uparrow}|^{2}/2}.$ (35)
Therefore, if |$\ket{\alpha_{\downarrow}}$ and |$\ket{\alpha_{\uparrow}}$ are identical, the spin is not measured and is not dephased. On the other hand, if |$\ket{\alpha_{\downarrow}}$ and |$\ket{\alpha_{\uparrow}}$ are orthogonal, then coherence is destroyed; this is the limit of projective measurement, in which maximal information may be acquired about the spin.
A.4.2 Measurement of the spin state
We now apply these general principles of quantum measurement to understand readout of a probed cavity QED system. Consider the model of an optical cavity coupled to a single “atom”, Fig. 8. As in Fig. 6, this atom is treated as a “lambda system” with spin ground states $\ket{\downarrow}$ and $\ket{\uparrow}$, which have transitions A1 or A2 to an excited state |$\ket{A}$, respectively. Coupling between the atom and cavity can change the cavity frequency by the dispersive shift ±$\pm\chi$, dependent on the spin state. The magnitude $\chi$ depends on the coupling between atom and cavity and their detuning is related to the cooperativity between them (desired to be high). A noticeable shift can occur, for instance, if the atom is tuned such that the A1 transition is resonant with the cavity, coupling between the A1 transition and cavity is strong compared to overall system loss, and the A2 transition is far off resonant (so coupling between this transition and the cavity is negligible). Such a system is well approximated by a limit of the Jaynes-Cummings Hamiltonian [74] $\hat{H}/\hbar\approx(\omega_{c}+\chi\hat{\sigma}_{z})\hat{a}^{\dagger}\hat{a}+\omega_{q}\hat{\sigma}_{z}/2$, where $\hat{a}$ is the cavity field operator, $\hat{\sigma}_{z}$ is the Pauli operator describing the atom’s spin state, $\omega_{c}$ is the bare cavity frequency, and $\omega_{q}$ is the spin transition frequency. Finally, readout requires nonunitary interaction with a measurement apparatus. The total loss rate of the cavity is described by the linewidth $\kappa$, which includes internal loss $\kappa_{\mathrm{int}}$ coming from intrinsic cavity loss of the loaded cavity, as well as external loss $\kappa_{\mathrm{ext}}$ due to external coupling.
Under a near-resonant drive we can use input-output theory to solve for the steady-state field inside the cavity. Driven far below saturation and in the limit where $\gamma\ll\kappa$, this becomes [67]
$\alpha_{\downarrow,\uparrow}(\delta)=\frac{\epsilon}{\sqrt{(\kappa/2)^{2}+(\delta\pm\chi)^{2}}},$ (36)
where $\delta$ is the detuning of the drive from the bare cavity frequency, the drive rate $\epsilon$ is proportional to the drive amplitude, and $\epsilon^{2}$ is proportional to drive power. (Relaxing the assumption of $\gamma\ll\kappa$ complicates Eq. 36, but retains this proportionality [75].) The response of Eq. 36 is plotted in Fig. 8b for a choice of $\chi=2\kappa$.
Readout works by inferring the spin state based on a measurement of |$\ket{\alpha_{\downarrow}}$ or |$\ket{\alpha_{\uparrow}}$. Consider a steady-state system where the cavity is driven for a readout window $\tau$ much longer than the transient dynamics of the field’s ring-up or -down time, such that $\tau\gg 1/\kappa$. The mean number of expected photon counts at a detector is $\bar{n}_{b}=\eta\kappa\tau|\alpha_{\downarrow}|^{2}$ if the cavity is in |$\ket{\alpha_{\downarrow}}$ and $\bar{n}_{d}=\eta\kappa\tau|\alpha_{\uparrow}|^{2}$ if the cavity is in |$\ket{\alpha_{\uparrow}}$. Here, $\kappa$ sets the overall rate at which photons leave the cavity and are, thus, incoherently lost. The measurement efficiency $\eta$ is again the overall fraction of these photons that are actually detected.
Notably, unlike the photoluminescence-based readout discussed elsewhere, here $\eta$ does not suffer from loss due to the emitter’s quantum efficiency or fraction of emission into the phonon sideband vs zero-phonon line. This is because we consider the limit of low enough excitation power ($p\ll p_{\mathrm{sat}}$), such that the scattered probe dominates signal and photoluminescence is negligible. Cyclicity becomes important only as probe power is increased so that excitation or photoluminescence appears as a concurrent effect.
In analogy with Eq. 6 in the main text, we model the difference in collected counts between the bright and dark distributions to be $\bar{n}_{b}-\bar{n}_{d}=\eta\kappa\tau(|\alpha_{\downarrow}|^{2}-|\alpha_{\uparrow}|^{2})$. We would, therefore, experimentally measure
$\bar{n}_{b}-\bar{n}_{d}=\mathcal{A}p,$ (37)
where $\mathcal{A}=\eta\kappa\tau\left(|\alpha_{\downarrow}|^{2}-|\alpha_{\uparrow}|^{2}\right)/p$ is a constant of proportionality that can be fit to experimental data in analogy to Fig. 4b. Notice that $\mathcal{A}$ is independent of drive power, because the intracavity photon number (either $|\alpha_{\downarrow}|^{2}$ or $|\alpha_{\uparrow}|^{2}$) scales as $\epsilon^{2}$, which is proportional to $p$.
In summary, the difference between the measured distributions associated with the spin being in $\ket{\uparrow}$ or $\ket{\downarrow}$ allows for determination of the spin state. These distributions grow further apart as integration time increases, drive power increases, and measurement efficiency increases. The limit where these distributions are sufficiently separated allows for a high-fidelity single-shot measurement. In practice, it is desirable to have a short integration time for faster readout and a low drive amplitude to avoid other dynamics, and so it remains important to have high efficiency.
A.4.3 Determining efficiency of a cavity QED system
We have seen that measurement of the probe both dephases the qubit and allows for determination of the spin state. As for the case of readout using photoluminescent emission (e.g. Eq. 8), we compare these effects to quantify the measurement efficiency $\eta$.
After projective measurement of a spin-dependent probe, the spin’s postmeasurement coherence decreases by Eq. 35. In a steady-state cavity that is continuously driven for duration $\tau$, the field in the cavity is continuously being lost and/or measured at rate $\kappa$, and continuously repopulated by the drive. We can understand the overall dephasing given by Eq. 35 to occur at some dephasing rate $\Gamma_{\phi}$, such that the spin’s phase coherence decays in magnitude by a factor of $e^{-\Gamma_{\phi}\tau}$ over time $\tau$. The dephasing rate
$\Gamma_{\phi}=\frac{1}{2}\left|\alpha_{\downarrow}-\alpha_{\uparrow}\right|^{2}\kappa$ (38)
is the overlap of the probe states times the loss rate of the cavity.
Like the collected signal in Eq. 37, the dephasing rate is proportional to drive power such that
$\Gamma_{\phi}=\mathcal{B}p,$ (39)
where $\mathcal{B}=\kappa\left|\alpha_{\downarrow}-\alpha_{\uparrow}\right|^{2}/(2p)$ is a proportionality constant that may be experimentally determined by the method described in Fig. 5. (Note that $\mathcal{B}$ is constant with power, because $\left|\alpha_{\downarrow}-\alpha_{\uparrow}\right|^{2}\propto p$.)
In analogy with Eq. 8, comparison between constants $\mathcal{A}$ and $\mathcal{B}$ yields the measurement efficiency
$\eta=\frac{\mathcal{A}}{2\mathcal{B}\tau}\times f.$ (40)
Here, the added factor $f=|\alpha_{\downarrow}-\alpha_{\uparrow}|^{2}/(|\alpha_{\downarrow}|^{2}-|\alpha_{\uparrow}|^{2})$ depends on the spin-dependent steady-state cavity field, which is a function of $\chi$ and $\kappa$. These parameters may experimentally determined by spectroscopic measurements of the cavity in order to determine $f$. In the limit $\kappa\ll\chi$, this simplifies further to $f\approx 1$ and $\eta\approx\mathcal{A}/(2\mathcal{B}\tau)$.
In summary, we have derived a general method to characterize the measurement efficiency of $\eta$ of a probed cavity QED system. Readout using this method relies on scattering of a probe tone off a dispersively shifted cavity, such that the spin state is inferred by measurement of the scattered field. While this method is fundamentally different than the spin readout via detection of photoluminescent emission discussed elsewhere in this work, both techniques are well described by the general theory of quantum measurement including measurement-induced dephasing. We emphasize that this section has direct analogy to well-established techniques in the readout of superconducting qubits in circuit QED architectures [58, 59, 67, 60, 61, 62].
Finally, we emphasize the applicability of this section to readout of color center spin qubits integrated into cavity QED systems, including recent experiments using SiV- centers [20, 19, 21, 22] or quantum dots [72]. In contrast to the SiV- (737 nm zero-phonon line), the shorter emission wavelength of the SnV- (620 nm) can present a technical challenge to nanofabrication, but dispersive readout of the SnV-’s spin state should work comparably well so long as similar cooperativity can be reached. Recent demonstrations of Purcell enhancement of an SnV- center integrated into a nanophotonic cavity suggest that the SnV- platform is close to this regime [33, 34, 38], and, thus, disperive readout of an SnV- spin is a promising avenue for future experiments.